### v1

# Generating Jackiw-Teitelboim Euclidean gravity from static three-dimensional Maxwell-Chern-Simons electromagnetism

Abstract
We consider pure three-dimensional Maxwell-Chern-Simons electrodynamics in the static limit. We show that this theory can be mapped onto a two-dimensional gravitational model in the first-order formalism of Riemannian manifolds with Euclidean signature, coupled to a real scalar field naturally interpreted as a dilaton. In this framework, the Newtonian and cosmological constants in two dimensions are fully determined by the electric charge. The solution to this gravitational model is found to be trivial: a constant dilaton field on a flat manifold. However, we introduce two distinct shifts of the spin-connection that transform the model into Jackiw-Teitelboim gravity. Specifically, we identify two additional solutions: a hyperbolic manifold with also a constant dilaton configuration; and a spherical manifold where, again, the dilaton assumes a constant, nonzero field configuration. In both non-flat cases, by employing the Gauss-Bonnet theorem in the specific cases of compact manifolds, we establish that the manifold’s radius is fixed by the cosmological constant (and, therefore, by the electric charge).
1 Introduction
The primary objective of this work is to establish a connection between three-dimensional electrodynamics and Jackiw-Teitelboim (JT) gravity models [1, 2, 3, 4, 5]. Accordingly, the main system of interest in this paper is two-dimensional gravity [1, 2, 6, 7, 8, 9, 10, 11, 3, 5, 12]. In particular, two-dimensional gravity has served as a robust framework for probing the quantization of gravity in higher dimensions. Its topological nature further motivates the investigation of topological field theories as candidates for quantum gravity in four dimensions (See, for instance, [8, 13, 14, 15] and references therein). Another compelling reason for studying gravity theories across different dimensions is that numerous physical systems can be effectively described by geometrodynamical models [16, 17, 18, 19, 20, 21, 22]. In such models, gravity (or geometrodynamics) provides an alternative framework for describing certain gauge theories.
Specifically, in the case of JT gravity, several potential applications and relevant results have been explored [23, 24, 25, 26, 27, 28, 4, 29, 5, 30, 31, 32]. It is particularly significant in black hole physics. While the Einstein-Hilbert model with cosmological constant in two dimensions simply defines a Plateau problem [33, 34], the JT model has more dynamical features due to the presence of the dilaton, being thus a good description for two-dimensional black holes and for a large group of near-extremal black holes in higher dimensions [5, 24, 35, 26]. Another area of great contribution from JT gravity is holography (see [36] and references therein). Typically, JT models have constant and negative curvature as solutions, stating an $AdS_{2}$ spacetime (two-dimensional Anti-de Sitter spacetime). In this manner, $AdS/CFT$ correspondence [37] can play an important role in this scenario in order to study quantum gravity in a string theory fashion. By analyzing such a lower-dimensional system, the structure of quantum gravity could emerge more evidently and provide some insights for higher dimensions. In addition, JT gravity is relevant in matrix models [38, 39], which can be related to Sachdev–Ye–Kitaev (SYK) systems [40, 41] to describe chaotic quantum systems in condensed matter physics. It turns out that the SYK model is deeply connected to black hole physics due to a duality to JT gravity via $AdS/CFT$ correspondence (see for example [42] and references therein).
The second system under consideration in this work is three-dimensional electrodynamics, which allows for the inclusion of the Chern-Simons (CS) topological term. Electromagnetic models of this type have applications in various branches of physics, extending beyond purely field theoretical investigations. Notably, they have been employed to describe aspects of topological insulators [43, 44], Weyl semi-metals and axion effects [45, 46], graphene layers [47, 48, 49, 50], and even three-dimensional gravity [51, 16, 52, 21, 22]. Furthermore, the consistency of CS theories at the quantum level is a highly valued feature. In fact, it has long been established that quantum CS theories are perturbatively finite [53, 54].
The approach we adopt to connect electrodynamics to gravity theories was previously investigated in two recent studies by the authors [55, 56]. In those works, we demonstrated that two-dimensional gravity can emerge from two-dimensional electrodynamics (including fermions) by defining a suitable mapping between the fields of both theories. This mapping originates from the isomorphism between the groups $U(1)$ and $SO(2)$. It is important to highlight that the mapping is rigorously defined in Euclidean spaces — i.e., electrodynamics formulated in $\mathbb{R}^{2}$ spacetime is mapped onto gravity in a two-dimensional manifold equipped with a Euclidean signature metric. As will become clear throughout the paper, this feature is highly beneficial in the present study.
At this point, it is also worth mentioning that the gravity models examined here are formulated within the first-order formalism (FOF) [57, 58, 59, 33, 34]. The FOF is a theoretical framework in which gauge symmetry is identified with local spacetime isometries, bringing gravity more closely related to standard gauge theories. Moreover, in this formalism, the fundamental fields are the vielbein (zweibein, in the specific case of two dimensions) and the spin-connection, rather than the metric tensor and affine connection used in the Palatini formalism [60], although the correspondence between both formalisms is well defined [61].
We begin this work by considering a three-dimensional electrodynamics action comprising Maxwell and CS terms, i.e., the Maxwell-Chern-Simons (MCS) theory. This theory is then dimensionally reduced by imposing the static limit, leading to a two-dimensional magnetic theory with an additional scalar field. After exploring the fundamental aspects of this system, we proceed to map it onto a gravity theory, following the formalism developed in [55, 56]. In this context, the gravitational constants (Newton’s constant and the cosmological constant) are entirely determined by the electric charge. Since the electric charge is known [62], it can be used to obtain explicit values for these gravitational constants. Interestingly, we find that Newton’s constant and the cosmological constant are significantly larger when compared to their four-dimensional experimental values [62]. Specifically, Newton’s constant is found to be $0.22eV^{-1}$, while the cosmological constant is valued as $0.09eV$.
Furthermore, we identify a Euclidean two-dimensional dilatonic gravity model whose solution is trivial: the dilaton field remains constant throughout the manifold, while the manifold itself is flat with minimal area. Nonetheless, we demonstrate that the model can be reformulated to describe JT gravity with two distinct nontrivial solutions. The first corresponds to a manifold with a negative constant curvature and a constant dilaton, i.e., the model admits hyperbolic spaces as solutions [63, 64]. The second non-flat solution describes a regular spherical manifold with, again, a constant dilaton configuration.
In the three cases (flat, hyperbolic, and spherical), the Gauss-Bonnet (GB) theorem can be employed to determine some extra properties of these manifolds. Particularly, if we consider compact surfaces, we can study: the flat torus; the $n$-holed torus (negative curvature); and the sphere (positive curvature). Except for the flat case, we are able to determine the length typical scale of such surfaces, finding it to be of the order of $\mu m$ (micrometers). This is possible since the GB theorem relates the area of surfaces to their Euler characteristics [65, 63], and because the area depends on the cosmological constant (explicitly calculated from the electric charge). In addition, it is crucial to emphasize that the CS term plays a fundamental role in achieving the JT model. In fact, within our framework, the JT term is directly proportional to the CS level of the theory.
This work is organized as follows. In Section 2, we define three-dimensional MCS electrodynamics, discuss its main features, and take its static limit. In Section 3 we map the static electrodynamics into a two-dimensional gravity theory in the FOF and explore its flat solution. Then, in Section 4 we connect this gravity theory with JT gravities and explore the main properties of these solutions. In Section 5, we compute, by topological arguments, the radius of some manifolds, specifically for the $n$-holed torii and the sphere. Finally, our conclusions and perspectives are displayed in Section 6.
2 Three-dimensional electrodynamics
In this section, we define three-dimensional MCS electrodynamics and, by imposing the static limit, we dimensionally reduce it to two spatial dimensions.
2.1 The Maxwell-Chern-Simons theory
The starting point is the most general action in three dimensions of pure electrodynamics which is polynomial in the fields and their derivatives and power counting renormalizable. The result is just the pure MCS action,
$S_{3qed}=\int\left(-\frac{1}{e^{2}}f\underline{\ast}f+\kappa af\right)\;,$ (2.1)
with $a=a_{\mu}dx^{\mu}$ being the gauge field, $f=\underline{d}a$ being the field strength while $\underline{\ast}$ characterizes the three-dimensional Hodge dual. The notation $\underline{d}=\partial_{\mu}dx^{\mu}$ stands for the three-dimensional exterior derivative. Greek indices run through $\{0,1,2\}$ with the 00th coordinate characterizing time. Moreover, we consider Minkowskian three-dimensional spacetime with negative signature as the stage of the MCS theory (2.1). We also remark that we are using natural units.
The first term in (2.1) is clearly the Maxwell action. Therefore, $e$ is the electric charge. The second term in (2.1) is the CS topological term and $\kappa$ is the CS level of the theory. The canonical dimensions of fields and parameters are given by $\left[a\right]=1$, $\left[e\right]=1/2$, and $\left[\kappa\right]=0$. The field equations can easily be derived from the action (2.1) by applying Hamilton’s principle to it. It provides
$\underline{d}\underline{\ast}f+\kappa e^{2}f=0\;,$ (2.2)
Moreover, Bianchi identities are also valid, namely,
$\underline{d}f=0\;.$ (2.3)
By performing a suitable and straightforward spacetime decomposition, one can easily infer that the field strength $f$ is related to the electromagnetic fields through
$f^{\mu\nu}\equiv\begin{pmatrix}0&-E^{1}&-E^{2}\\ E^{1}&0&B\\ E^{2}&-B&0\end{pmatrix}\;.$ (2.4)
Such field strength describes an electrodynamics governed by a two-dimensional electric vector field and a pseudo-scalar magnetic field, as expected. The equations of motion can be decoupled in the usual way to generate massive wave equations,
$\left(\Box-m^{2}\right){\bf E}$ $=$ $0\;,$
$\left(\Box-m^{2}\right)B$ $=$ $0\;,$ (2.5)
with $\Box=\partial_{t}^{2}-\nabla^{2}$ being the d’Alambertian operator, ${\bf E}=(E^{1},E^{2})$, and $m=\kappa e^{2}$ being the CS topological mass. At the classical level, these waves decay exponentially on a characteristic length scale of the order of $m^{-1}$. This is the usual result of MCS electromagnetism.
Finally, it is convenient to recall that electrodynamics is known to be a gauge theory invariant under $U(1)$ gauge transformations,
$\delta a=\underline{d}\alpha\;,$ (2.6)
with $\alpha$ being a local parameter. In fact, electrodynamics can be fundamentally defined as a $U(1)$ gauge theory [66]. It turns out that the $U(1)$ symmetry is the key for the mapping connecting this theory to gravity.
2.2 Dimensional reduction - The static limit
Once the MCS is defined, dimensional reduction of the model can be performed. We choose to reduce the time coordinate by considering only static systems. This is the simplest way to reduce the time dimension. Therefore, decomposing space and time of the exterior derivative, the gauge field, and field strength, we find
$\underline{d}$ $=$ $\partial_{t}dt+d\;,$
$a$ $=$ $\Phi dt+A\;,$
$f$ $=$ $\left(d\Phi-\partial_{t}A\right)dt+F\;,$
¯$\underline{\ast}f$ $=$ $-\ast\left(d\Phi-\partial_{t}A\right)+\ast Fdt\;,$ (2.7)
with $A=A_{i}dx^{i}$ being the two-dimensional gauge field, $\Phi$ being the electric potential, and $F=dA$ being the two-dimensional electromagnetic field strength. Moreover, $\ast$ is the Hodge dual in the two-dimensional Euclidean space while $d=\partial_{i}dx^{i}$ is the purely spatial exterior derivative. The Latin indices $i,j,k\ldots$ run through $\{1,2\}$, and describe the Euclidean spatial sector. Thus, the Lagrangian terms decompose accordingly,
$f\underline{\ast}f$ $=$ $\left[\left(d\Phi-\partial_{t}A\right)\ast\left(d\Phi-\partial^{t}A\right)+F\ast F\right]dt\;,$
$af$ $=$ $\left(\Phi F+Ad\Phi-A\partial_{t}A\right)dt\;.$ (2.8)
The corresponding decomposition of the action (2.1) is just
[…]
$S_{3stqed}=\int\left[-\frac{1}{e^{2}}\left(d\Phi\ast d\Phi+F\ast F\right)+\kappa\left(\Phi F+Ad\Phi\right)\right]\;,$ (2.10)
where integration is taken over the two-dimensional Euclidean space. Moreover, since $Ad\Phi=\Phi F-d(\Phi A)$, we are allowed to rewrite action (2.10) as
$S_{3stqed}=\int\left\{-\frac{1}{e^{2}}\left(d\Phi\ast d\Phi+F\ast F\right)+\kappa\left[2\Phi F-d\left(\Phi A\right)\right]\right\}\;,$ (2.11)
simplifying thus the term $Ad\Phi$ to a surface term. Of course, the surface term can be relevant in systems with non-trivial boundaries, like topological insulators or similar systems.
The action (2.11) is just a sort of three-dimensional static electromagnetism that can be viewed in different ways. For instance, it can be interpreted as electrodynamics in two Euclidean dimensions coupled to a scalar field; equivalently, still under the two-dimensional point of view, one can interpret the static model as a scalar field $\Phi$ interacting with a pseudo-scalar magnetic field through the reduced field strength
$F^{ij}\equiv\begin{pmatrix}0&B\\ -B&0\end{pmatrix}\;.$ (2.12)
It is worth mentioning that the present situation is essentially different from the previous works [55, 56] where the two-dimensional electrodynamics considered was purely electrical. It is possible to achieve a purely electrical system by reducing the third dimension ($\mu=2$) instead of the time coordinate. However, the resulting electrical model would still be living in Minkowskian spacetime. Such a model could not be mapped in a gravity model in an easy way as in the Euclidean case. This is the main reason why we dimensionally reduce time instead of a spatial dimension. See [55, 56] for further details.
The static field equations can be derived directly from the field equations (2.2) or from the minimization of the action (2.11). The field equations of the scalar and vector fields read
$d\ast d\Phi+mF$ $=$ $0\;.$
$d\ast F+md\Phi$ $=$ $0\;.$ (2.13)
These equations can be decoupled into
$\left(\nabla^{2}-m^{2}\right)\Phi$ $=$ $C\;,$
$\left(\nabla^{2}-m^{2}\right)A$ $=$ $G\;,$ (2.14)
with $G$ and $C=C_{i}dx^{i}$ being integration constants that can be chosen to vanish, with no loss of generality. These are massive Laplace equations whose solutions, in polar coordinates, are given in terms of the modified Bessel functions of the first and second kinds depending on $mr$ (with $r$ being the radial coordinate), and periodic functions in the angular coordinate.
It is also useful, for future purposes, to infer that, after the dimensional reduction, the gauge transformations (2.6) decouple to
$\Phi^{\prime}$ $=$ $\Phi\;,$
$A^{\prime}$ $=$ $A+d\alpha\;,$ (2.15)
i.e., the scalar field becomes invariant under gauge transformations while the magnetic vector potential remains as a gauge field.
As an extra and final step, we rewrite the action (2.11) with the help of an auxiliary field (See [17, 55, 56]),
$S_{3qed}=\int\left\{-\frac{1}{e^{2}}\left(d\Phi\ast d\Phi+\Theta\ast F-\frac{1}{2}\Theta\ast\Theta\right)+\kappa\left[2\Phi F-d\left(\Phi A\right)\right]\right\}\;.$ (2.16)
The field equation of the auxiliary field $\Theta$ trivially reads $\Theta=F$, which implies in the action (2.11) again. Of course, the introduction of auxiliary fields in this way is also valid at the quantum level [67].
To end up this Section, for completeness, and usefulness, the canonical dimensions, and form ranks of the fields and parameters appearing in the action (2.16) are displayed in Table 1.
Fields $A$ $\Theta$ $\Phi$ $e$ $\kappa$
Dimension $1$ $2$ $1$ $1/2$ 00
Rank $1$ $2$ 00 00 00
Table 1: Canonical dimensions and form ranks of the electromagnetic fields and parameters.
3 Achieving gravity and the flat solution
In this section we map the static electromagnetic action (2.16) into a two-dimensional gravity theory in manifolds of Euclidean signature. The details and formal aspects of the mapping can be found in [55, 56]. In those papers, extra auxiliary non-physical fields are introduced in order to make the mapping well defined. In here, we omit such particularities because the same method applies in the present case. In a nutshell, the mapping is based on the isomorphism between the gauge group $U(1)$ of electrodynamics and the $SO(2)$ rotations of Euclidean two-dimensional gravity. The map between the fields of electrodynamics and gravity is [^1] The factor $e^{4}$ is different from the pure two-dimensional case [55, 56] because the canonical dimension of the electric charge is different. In fact, if $D$ is spacetime dimension, the canonical dimension of the electric charge is $[e]=(4-D)/2$.,
[…]
with $e^{a}=e^{a}_{i}dX^{i}$ being the zweibein 1-form, with vanishing mass dimension $[e]=0$, while $\omega^{ab}=\omega^{ab}_{i}dX^{i}$ is the spin-connection, with mas dimension given by $[\omega]=1$. The point $x\in\mathbb{R}^{2}$ is a point in the Euclidean space where action (2.16) is defined and $X\in M^{2}$ is a point in the two-dimensional manifold $M^{2}$ where gravity will emerge. The indices $a,b,c,\ldots,h\in\{1,2\}$ are associated with the tangent space of $M^{2}$, essentially a local inertial frame, while $i,j,k,\ldots\in\{1,2\}$ refers to a general coordinate system of the manifold $M^{2}$. The essence of the mapping is the identification of the fields of three-dimensional static electromagnetism with the geometric fields of the manifold, i.e., the dynamics of the electromagnetic fields are absorbed by the manifold.
Performing the mapping (3.1) in the action (2.16), one straightforwardly finds
$S_{2grav}=\int\left\{-\frac{1}{e^{2}}\left(d\Phi\ast d\Phi+2e^{4}\epsilon_{ab}R^{ab}-e^{8}\epsilon_{ab}e^{a}e^{b}\right)+\kappa\left[2\Phi\epsilon_{ab}R^{ab}-\epsilon_{ab}d\left(\Phi\omega^{ab}\right)\right]\right\}\;,$ (3.2)
with $R^{ab}=d\omega^{ab}$ being the curvature 2-form and $\epsilon_{ab}$ being the Levi-Civita symbol. Therefore, inferring the parameter identifications [^2] Obviously, $[G]=-1$ and $[\Lambda]=1$.,
$G$ $=$ $\frac{1}{16\pi e^{2}}\;,$
$\Lambda$ $=$ $e^{2}\;,$ (3.3)
we find the gravity model,
$S_{2grav}=\int\left\{-\frac{1}{\Lambda}d\Phi\ast d\Phi-\frac{1}{8\pi G}\epsilon_{ab}\left(R^{ab}-\frac{\Lambda^{2}}{2}e^{a}e^{b}\right)+\kappa\left[2\Phi\epsilon_{ab}R^{ab}-\epsilon_{ab}d\left(\Phi\omega^{ab}\right)\right]\right\}\;.$ (3.4)
This is a two-dimensional gravity action, in the first-order formalism, coupled with a scalar field, naturally identified as the dilaton. The constants $G$ and $\Lambda$ are recognized as the two-dimensional Newtonian and cosmological constants, respectively. Remarkably, since we know the value of the electric charge, we can compute the numerical values of the gravitational constants through expressions (3.3). Indeed,
$G$$\approx$ $0.22eV^{-1}\;,$
$\Lambda$$\approx$ $0.09eV\;.$ (3.5)
Hence, $G$ and $\Lambda$ are very large if compared with their four-dimensional physical values [62], namely $G\approx 6.7\times 10^{-57}eV^{-2}$ and $\Lambda\approx 2.1\times 10^{-33}eV$. Moreover, from (3.3), these constants are constrained by
$G\Lambda=\frac{1}{16\pi}\;.$ (3.6)
The first term of the action (3.4) is the standard kinematic term for a real scalar field, up to a constant factor. The second and third pieces are the typical Einstein-Hilbert (EH) and cosmological constant terms. The EH action, in two dimensions, is known to be of topological nature, a property evidenced by the fact that it is a total derivative. In fact, in two dimensions, the EH term coincides with the GB term. See Section 5.
Continuing the analysis of the action (3.4), the next term is the dilaton-curvature coupling, emerging from the CS term. The last term is just a total derivative. This term, just like the EH one, could only contribute to the field equations at the boundary if this boundary is non-trivial.
Clearly, the action (3.4) is, up to boundary terms, manifestly gauge invariant under $SO(2)$ rotations in tangent space,
$\delta\Phi$ $=$ $0\;,$
$\delta e^{a}$ $=$ $\alpha^{a}_{\phantom{a}b}e^{b}\;,$
$\delta\omega^{ab}$ $=$ $d\alpha^{ab}\;.$ (3.7)
Thus, the action (3.4) is a genuine Euclidean two-dimensional gravity, in the FOF, coupled to a scalar field. We point out that transformations (3.7) are directly obtained from (2.15), (3.1) and the identification $\alpha(x)=\epsilon_{ab}\alpha^{ab}(X)$, see [55, 56]. We also stress that action (3.4) displays diffeomorphism symmetry, which is implicit in form notation.
In the next section, we will show that this action can be transmuted in JT-type gravities. For now, let us just exploit the dynamics of the action (3.4). The field equation of $\Phi$ is just
$d\ast d\Phi+\Lambda\kappa\epsilon_{ab}R^{ab}=0\;,$ (3.8)
suggesting that the scalar field carries some dynamics. It also suggests that the scalar field affects non-trivially the manifold geometry. However, the spin-connection equation gives
$d\Phi=0\;,$ (3.9)
So, the dilaton solution is simply an arbitrary constant field configuration, possibly zero. Furthermore, combining equations (3.8) and (3.9) one easily finds that the manifold is locally flat,
$R^{ab}=0\;.$ (3.10)
It remains to derive the equation of the zweibein. The only non-trivial term remaining in the action (3.4) that depends on the zweibein is the area term. Then, the final equation is simply the minimization of the area of the manifold (Plateau problem), just like two-dimensional Mardones-Zanelli gravity [33, 34]. The solution here is thus a flat space with minimal area. Of course, the specific manifold depends on the boundary conditions. Typical non-compact examples are planes, cylinders, and disks. In the case of compact manifolds, the flat torus is the standard solution. See Figure 1, in Section 5, for examples. In fact, we will return to the discussion of these surfaces in Section 5.
4 Euclidean Jackiw-Teitelboim gravity
In this section, we demonstrate that the action (3.4) can be related to the JT model [1, 2, 5], resulting in two types of manifolds: hyperbolic and spherical.
4.1 Hyperbolic solution
The first way to obtain a JT gravity is by carrying out the curvature shift
$R^{ab}\longrightarrow R^{ab}+\frac{\Lambda^{2}}{2}e^{a}e^{b}\;,$ (4.1)
enforcing the JT usual term to directly appear in the gravity action (3.4). However, before we perform this shift in the proper way, let us define it a bit more rigorously. In fact, the shift (4.1) is actually generated by a shift in the spin-connection, namely
$\omega^{ab}\longrightarrow\omega^{ab}+\frac{\Lambda^{2}}{2}X^{ab}\;,$ (4.2)
with the local condition
$dX^{ab}=e^{a}e^{b}\;,$ (4.3)
meaning that the field $X$ must be a non-conservative field with a canonical dimension of $[X]=-1$. For further reference, we name condition (4.3) shift constraint. Thus, once the zweibein $e$ describing the manifold is chosen (note that there are infinite possibilities due to the $SO(2)$ gauge symmetry and diffeomorphisms), the field $X$ can be determined. Integration of equation (4.3), and resorting to Stokes’ theorem, gives the global condition
$\oint_{C}X^{ab}=\epsilon^{ab}\int_{S}d^{2}x\;.$ (4.4)
Thus, the circulation of $X$ is essentially determined by the area $S$ inside the closed curve of circulation $C=\partial S$. Typically, $S$ is of free choice; for instance, one can set $S\equiv M^{2}$, if $\partial M^{2}\neq 0$.
Another point of importance concerning equation (4.3) is that it is not gauge covariant unless $dX$ is covariant, which is actually not true. In fact, the shift (4.2) suggests that $X$ is a gauge invariant field. As a consequence, gauge symmetry is broken by the constraint (4.3); it may be seen, thus, as a gauge fixing for the model. In fact, together with the shift (4.2), it can be incorporated into the action (3.4) with a suitable 0-form Lagrange multiplier [^3] The Lagrange multiplier carry mass dimension 3, $[\lambda]=3$. $\lambda_{ab}$ by means of the addition of the action
$S_{\lambda}=\int\lambda_{ab}\left(dX^{ab}-e^{a}e^{b}\right)\;.$ (4.5)
This term clearly breaks gauge symmetry, assuming that $\lambda$ is either gauge invariant or covariant.
Therefore, by employing the shift (4.2) in the action (3.4) and the shift constraint via the action (4.5), we have the full action
$S_{JT\lambda 1}$ $=$ $\int\left\{-\frac{1}{\Lambda}d\Phi\ast d\Phi-\frac{1}{8\pi G}\epsilon_{ab}\left[R^{ab}+\frac{\Lambda^{2}}{2}\left(dX^{ab}-e^{a}e^{b}\right)\right]+2\kappa\Phi\epsilon_{ab}\left(R^{ab}+\frac{\Lambda^{2}}{2}dX^{ab}\right)+\right.$ (4.6)
$+$ $\left.\lambda_{ab}\left(dX^{ab}-e^{a}e^{b}\right)+\kappa\epsilon_{ab}d\left(\Phi\omega^{ab}+\frac{\Lambda^{2}}{2}\Phi X^{ab}\right)\right\}\;.$
Let us derive the field equations. The equation for the Lagrange multiplier naturally imposes the shift constraint (4.3). The equation of the spin-connection remains the same, equation (3.9). Thus, $\Phi$ is again undetermined. It is constant, but not necessarily zero. However, nothing forbids us to choose $\Phi=0$. The equation for $X$, provides (already employing (3.9))
$d\lambda^{ab}=0\;.$ (4.7)
implying that $\lambda$ has a constant configuration.
Computing the field equation for $\Phi$, already applying (3.9) and the equation for $\lambda$ (the shift constraint), provides the equation for hyperbolic manifolds,
$R^{ab}+\frac{\Lambda^{2}}{2}e^{a}e^{b}=0\;,$ (4.8)
which is just JT solution for negative constant curvature in Euclidean spaces.
Now, the zweibein field equation, already neglecting the kinematical term due to (3.9), gives
$\lambda_{ab}=\frac{\Lambda^{2}}{16\pi G}\epsilon_{ab}\;.$ (4.9)
Remarkably, due to the constraint (3.6), $\lambda$ can be written in terms of the cosmological constant:
$\lambda_{ab}=\Lambda^{3}\epsilon_{ab}\;.$ (4.10)
This is actually quite an interesting result because a Lagrange multiplier works as a force obliging the system to obey the constraint. It turns out that the cosmological constant is that force. In essence, the cosmological constant is the force that bends the manifold.
The solution we found so far is just a dilatonic trivial vacuum and a manifold with negative constant curvature, i.e., a hyperbolic surface. In two-dimensional Euclidean spaces, such manifolds, if compact, are the connected $n$-holed torii [63, 64] - See Figure 3 and Section 5 for further discussions.
4.2 Spherical solution
The second way to achieve a JT solution is to perform the same shift (4.1), but with a different sign at the cosmological constant factor,
$R^{ab}\longrightarrow R^{ab}-\frac{\Lambda^{2}}{2}e^{a}e^{b}\;.$ (4.11)
Just like the previous case, the shift (4.11) demands a more fundamental shift, here described by
$\omega^{ab}\longrightarrow\omega^{ab}-\frac{\Lambda^{2}}{2}X^{ab}\;.$ (4.12)
Therefore, the shift constraint (4.3) (or equivalently, (4.4)) remains valid. In fact, the whole discussion about the shift constraint in the hyperbolic case remains valid here. Thence, performing the shift (4.12) in the gravity action (3.4) and adding the gauge fixing term (4.5) to the action results in the full action
$S_{JT\lambda 2}$ $=$ $\int\left\{-\frac{1}{\Lambda}d\Phi\ast d\Phi-\frac{1}{8\pi G}\epsilon_{ab}\left[R^{ab}-\frac{\Lambda^{2}}{2}\left(dX^{ab}+e^{a}e^{b}\right)\right]+2\kappa\Phi\epsilon_{ab}\left(R^{ab}-\frac{\Lambda^{2}}{2}dX^{ab}\right)+\right.$ (4.13)
$+$ $\left.\lambda_{ab}\left(dX^{ab}-e^{a}e^{b}\right)+\kappa\epsilon_{ab}d\left(\Phi\omega^{ab}-\frac{\Lambda^{2}}{2}\Phi X^{ab}\right)\right\}\;.$
Let us compute the field equations. Equation (3.9) remains valid, stating that $\Phi$ is a constant field again. The equation of $\lambda$ simply enforces the shift constraint (4.3). Again, the equation of the field $X$ gives that $d\lambda=0$, due to (3.9). Thus, $\lambda$ is also a constant field. The field equation of the dilaton, already employing the equations of $\omega$ and $\lambda$, gives the JT solution for a spherical manifold (constant positive curvature),
$R^{ab}-\frac{\Lambda^{2}}{2}e^{a}e^{b}=0\;.$ (4.14)
Finally, the equation for the zweibein again provides a fixed value for the Lagrange multiplier, see Equations (4.9) and (4.10). And the dilaton remains undetermined, with the freedom of setting it to vanish.
4.3 Extra remarks about the shift constraint
To finalize this Section, we remark that the shift constraint (4.3) can also be visualized as the imposition of a background. Therefore, the gauge fixing is actually a choice of an appropriate background. The only difference between the hyperbolic and the spherical cases is just the relative sign of the background. Essentially, the shifts (4.2) and (4.12), together with the constraint (4.3), impose the background to be hyperbolic or spherical. Moreover, since in two dimensions the EH term is topological, there is no room to generate Einstein-like equations in order to provide dynamics around the background.
In fact, the shifts (4.2) and (4.12) are simply stating that ±$\sim\pm X$ is the spin-connection background in each case. The shift constraint (4.3) states that $X$ must be the field that generates the specific (hyperbolic or spherical) background curvature. Moreover, it is known that the spin connection carries a specific term depending on the zweibein. Therefore, it is only natural that the shift constraint is written in terms of the zweibein. This is even evidenced by the fact that torsion is zero in our solutions.
One important piece of evidence that the model is not changed and the shift constraint, together with the shifts, is actually a background choice is the computation of the saddle point of actions (3.4), (4.6), and (4.13). They all converge to the same value, namely
$S_{2grav}\Big{|}_{on-shell}=S_{JT\lambda 1}\Big{|}_{on-shell}=S_{JT\lambda 2}\Big{|}_{on-shell}=\frac{\Lambda^{2}}{16\pi G}\mathcal{A}\;,$ (4.15)
where $\mathcal{A}$ is the area of the manifold (or, at least, a specific piece of the manifold, if the manifold is not completely integrable).
It is also interesting to notice that the JT term could emerge directly from the mapping by considering an electromagnetic background field. Right from the action (2.16), one could consider the shift $F\longrightarrow F\pm\Theta/2$. Or, equivalently, a shift on the electromagnetic field $a\longrightarrow a\pm M$ with $M$ being fixed by $dM=\Theta/2$. In this way, the target gravity action would already contain the usual JT term $\Phi(R\pm\Lambda^{2}/2)$. Therefore, the shift we proposed is equivalent to the existence of a background electromagnetic field in the original MCS theory.
Another comment is that the shift constraint (4.3) induces a relation between the spin-connection and the field $X$. This relation is actually generated by the cohomology of the nilpotent exterior derivative $d$, the well-known de Rham’s cohomology [65]. In fact, the constraint (4.3) and the curvature equations (4.8) (negative sign) and (4.14) (positive sign) can be combined to deduce the on-shell relation
$\omega^{ab}=\pm\frac{\Lambda^{2}}{2}X^{ab}+Y^{ab}\;,$ (4.16)
with $Y$ being a closed 1-form with dimension $[Y]=1$,
$dY^{ab}=0\;.$ (4.17)
Moreover, since $X$ is gauge invariant, $Y$ must transform as a connection in order to preserve the gauge structure of relation (4.16), namely
$\delta Y^{ab}=d\alpha^{ab}\;.$ (4.18)
Therefore, the gauge invariant field $X$ is essentially a difference between two connections, $X=\mp 2(\omega-Y)/\Lambda^{2}$, i.e., the difference between the spin-connection and an arbitrary closed 1-form.
Condition (4.17) is actually a traditional cohomology problem whose solution is simply given by its non-trivial and trivial parts in de Rham’s cohomology, namely
$Y^{ab}=Z^{ab}+dN^{ab}\;,$ (4.19)
with $Z$ being a closed non-exact 1-form with dimension $[Z]=1$,
$dZ^{ab}=0\;\Big{|}\;Z^{ab}\neq d(\mathrm{something})\;,$ (4.20)
while $N$ is an arbitrary 0-form (consequently, it cannot be written as an exact quantity) with dimension $[N]=0$. The gauge transformations of $Z$ and $N$ can be inferred from (4.18) and (4.19). Clearly, $Z$ can be chosen to be a gauge connection while $N$ is set to be gauge invariant [^4] Alternatively, there is also the possibility to choose $Z$ to be a gauge invariant field while $N$ would transform in a non usual way by $\delta N=\alpha$, i.e., $N$ would translate under $SO(2)$ gauge transformations. Such transformation indicates that $N$ could be interpreted as a field of angular nature. In this case, the relation (4.19) could be seen as a gauge transformation of the field $Y$ while $N$ would be the gauge parameter.,
$\delta Z^{ab}$ $=$ $d\alpha^{ab}\;,$
$\delta N^{ab}$ $=$ $0\;.$ (4.21)
The determination of the spin-connection works as follows. Once the zweibein describing the manifold is chosen, one can find any $X$ satisfying the shift constraint (4.3), a closed non-exact field $Z$, satisfying (4.20), and any gauge invariant 0-form $N$. Then, from relations (4.16) and (4.19), the spin-connection is completely determined. The same reasoning works to find $X$ by solving equation (4.8) (or (4.14)) for the spin-connection. In essence, the spin-connection and the field $X$ differ only by any non-trivial field $Z$ plus an exact field. Moreover, $Z$ and $N$ take no part in the solutions we found for JT gravity, see Equations (4.8) and (4.14). Alternatively, $X$ and $\omega$ can be deduced independently from equations (4.3) and (4.8) (or (4.14)), respectively. Then, $Z$ and $N$ can be inferred by comparing $X$ and $\omega$ via relation (4.16).
It is also important to stress that $Y$ must be non-vanishing because of the gauge transformations of $\omega$, $Y$ (see (3.7) and (4.17)), and the gauge invariant field $X$. Otherwise, relation (4.16) would not be gauge covariant. Thus, if $Z\neq 0$, then $dN$ can be set to vanish. On the other hand, since $N$ is gauge invariant, $Z$ cannot be set to zero. Therefore, from now on, we are allowed to fix $N=0$, with no loss of generality. As an alternative, if we do not care about gauge covariance (which is a valid choice since the presence of $X$ in the action already breaks gauge symmetry), we can set $Z=N=0$ and work with a gauge invariant spin-connection. However, we will not follow that path.
Now that we have all necessary fields in the gravity model, we display the dimensions and form ranks of all fields and parameters in Table 2.
Fields $e^{a}$ $\omega^{ab}$ $\Phi$ $X$ $Y$ $Z$ $N$ $\lambda$ $G$ $\Lambda$ $\kappa$
Dimension 00 $1$ $1$ $-1$ $1$ $1$ 00 $3$ $-1$ $1$ 00
Rank $1$ $1$ 00 $1$ $1$ $1$ 00 00 00 00 00
Table 2: Canonical dimensions and form ranks of the fields and parameters in the gravity theory.
The set of fundamental fields living in $M^{2}$ is thus =$\mathcal{F}=\{e,\omega,X,Z,\Phi,\lambda\}$, interlinked by the field equations
$d\Phi$ $=$ $dZ^{ab}\;\;=0\;,$
$\lambda_{ab}$ $=$ $\Lambda^{3}\epsilon_{ab}\;,$
$R^{ab}$ $=$ ±$\pm\frac{\Lambda^{2}}{2}e^{a}e^{b}\;,$
$\omega^{ab}$ $=$ ±$\pm\frac{\Lambda^{2}}{2}X^{ab}+Z^{ab}\;.$ (4.22)
The third equation in (4.22) establishes that, among $\omega$, $X$, and $Z$, we can neglect one of them in favor of the other two. The space of fundamental fields is then reduced, for instance, to $\mathcal{F}_{o}=\{e,X,Z,\Phi,\lambda\}$, where we have chosen to omit the spin-connection.
We also remark that in all three cases (flat, hyperbolic, and spherical), the action has no torsional terms, with the torsion 2-form being defined as the covariant derivative of the zweibein, $T^{a}=De^{a}=de^{a}+\omega^{a}_{\phantom{a}b}e^{b}$. Nevertheless, torsion could be generated depending on the gravity-matter coupling. In dimensions greater than two, torsion and curvature are related by the Bianchi identity $DT\sim Re$ [33, 34]. Thence, $T$ could be inferred from this relation. However, in two dimensions, this Bianchi identity vanishes identically, leaving torsion to be completely free [^5] The same effect occurs in pure the Mardones-Zannelli gravity [33, 34].; it depends solely on the choices we make in the solutions for $e$ and $\omega$. Therefore, one can simply set torsion to a fixed value and work one extra equation together with (4.22). For instance, choosing torsion to be zero and using the last relation in Equations (4.22), it is found that $Z$ (or $X$) is no longer independent,
$Z^{a}_{\phantom{a}b}e^{b}=-de^{a}\mp\frac{\Lambda^{2}}{2}X^{a}_{\phantom{a}b}e^{b}\;,$ (4.23)
The set of independent fields can be thus even reduced to $\mathcal{F}_{0}=\{e,Z,\Phi,\lambda\}$.
5 Gauss-Bonnet theorem: computing sizes
The GB theorem [^6] In the discussion provided here of the GB theorem, we attain ourselves to two-dimensional Riemannian manifolds. For generalizations to any dimension, see [68, 69]. [63, 65] is a powerful tool in the study of manifolds. The standard GB theorem for a connected and compact manifold, say $M^{2}$, states that [63, 65],
$\int_{M^{2}}\epsilon_{ab}R^{ab}=4\pi\chi\;,$ (5.1)
with $\chi$ being the Euler characteristic of the manifold. It is known that the GB theorem (5.1) generalizes for complete, finitely connected, and non-compact manifolds [70, 71, 72, 73] if the integral is absolutely integrable and the total area is finite. Nevertheless, we mostly confine ourselves to the compact surface cases.
Formula (5.1) can be employed to select more accurately the possible manifolds of our solutions. And, in some cases, it can be employed to determine their sizes, according to their Euler characteristics.
5.1 Flat surfaces
Solution encountered in Section 3 is a locally flat manifold of minimal area. Considering the GB theorem for compact surfaces of vanishing curvature, GB theorem in the form (5.1) restricts the pool of solutions to manifolds with zero Euler characteristic. It turns out that the manifold satisfying all requirements of the GB theorem is the torus with only one handle [63].
It turns out that formula (5.1) can also be applied for non-compact surfaces such as the cylinder [63, 74]. In fact, any manifold topologically equivalent to a cylinder or a torus, while maintaining the flatness of the surface, can be considered as a solution. For example, a disk with a hole in it satisfies the criteria. Observe that the cylinder and the holed disk are topologically equivalent; both are obtained from the continuous deformation of a sphere with two holes.
Unfortunately, the size of the surfaces with vanishing curvature cannot be computed from the GB theorem. Nevertheless, the theorem can be used to select the possible solutions to be in the set of surfaces with $\chi=0$. Some examples of flat surfaces with vanishing Euler characteristics are displayed in Figure 1.
Figure 1: Flat surfaces with vanishing Euler characteristics. From left to right: The cylinder; the flat disk with a hole; the flat torus. Among these manifolds, only the flat torus is compact.
5.2 Hyperbolic surfaces
The solution found in Section 4.1 is a manifold of negative constant curvature, i.e., any hyperbolic surface [63, 64]. Typical non-compact examples are the tractroid (pseudo-sphere) and the catenoid (See Figure (2)), which is homeomorphic to the cylinder, and thus has a vanishing Euler characteristic as well [63]. Therefore, we cannot employ the GB theorem (5.1) to compute its size. For compact hyperbolic surfaces, the simplest manifolds with negative curvature are essentially those which are equivalent to the $n$ linearly connected torii ($n$-holed torus), where $n$ is the number of handles (related to its genus) and the corresponding Euler characteristic is given by $\chi=2-2n$, with $n>1$ (The case $n=1$ is the simple torus with only one handle and vanishing curvature - see the previous Section) [63, 64]. Clearly, for hyperbolic surfaces, $\chi$ must be negative. Of course, the torii can be non-linearly glued, but the resulting genus would be bigger than $n$, as we will approach in a following example. Thus, putting the solution (4.8) in the GB theorem (5.1), we find [^7] We remark that it is commonly known the validity of the formula $\mathcal{A}_{h}=-2\pi\chi\;,$ (5.2) for the area of compact surfaces [63]. At a first look, this formula seems to not be compatible with formula (5.4). Nevertheless, this formula is deduced for surfaces with normalized curvature, which is not the case here due to the presence of the cosmological constant.
$\mathcal{A}_{h}=\frac{8\pi(n-1)}{\Lambda^{2}}\;,$ (5.3)
with $\mathcal{A}_{h}$ being the area of the $n$-holed torus. Substituting the value of the cosmological constant computed in expression (3.5) in formula (5.3), gives
$\mathcal{A}_{h}=3,102.8(n-1)eV^{-2}\;,$ (5.4)
which, in SI units, corresponds to
$\mathcal{A}_{h}=120.42(n-1)\mathrm{\mu m}^{2}\;.$ (5.5)
Figure 2: Non-compact hyperbolic manifolds. From left to right: The tractroid; and the catenoid. These manifolds have vanishing Euler characteristics.
Let us consider a first example, for instance (See Figure 3), $n$ identical torii, linearly connected, with two equal radii $2b_{h}$ and $b_{h}$, and making the interface between each torii be a circle of radius $c<<b_{h}$ (In this way, we characterized the manifold by only one parameter). This is the well known $n$-holed torus. We can determine the radius $b_{h}$ from formula (5.5):
$b_{h}=1.23\sqrt{\frac{(n-1)}{n}}\mu m\;.$ (5.6)
The minimum possibility is thus $n=2$, providing $b_{h}=0.87\mu m$. As $n\longrightarrow\infty$, the area stabilizes at the maximum value of $b_{h}=1.23\mu m$. We conclude that, although the area diverges with $n$, the radii of the identical $n$-holed torus do not.
Figure 3: The $n$-holed torus. A compact hyperbolical surface with non-vanishing Euler characteristics ($\chi=2-2n$). To compute the size of this manifold in an easy way, the intersections of the torii are made very small.
Just to illustrate another case, where the genus changes, we consider now the same $n$-holed torus of the previous example. However, we glue the first and last torii together, in a kind of ring made by $n$ torii, see Figure 4. In this case, the genus increases by one, $\chi=2-2(n+1)=-2n$. Thence,
$\mathcal{A}_{h}=\frac{8\pi n}{\Lambda^{2}}\;,$ (5.7)
implying that $b_{h}$ no longer depends on $n$ and is fixed by the value $b_{h}=1.23\mu m$.
Figure 4: The closed $n$-holed torus. They are linearly connected and the torii of the edges are also connected, forming a ring of torii. The big hole in the middle increases its Euler characteristics by one ($\chi=-2n$). Just like the open linearly connected $n$-holed torus, the intersections of the torii are made very small in order to compute the size of this manifold in an easy way.
These results are quite remarkable because it essentially says that the sizes of some surfaces in the present gravity model are determined ultimately by the electric charge in the original model, i.e., the manifold scale in one theory is related to the electric charge in another theory.
5.3 Spherical surfaces
In Section 4.2, another sort of solution was derived. Essentially, manifolds with constant positive curvature (see (4.14)). Traditional non-compact examples are spherical caps and curved disks. For compact solutions, we have the usual sphere itself.
Figure 5: The sphere. A compact manifold with positive curvature and non-trivial Euler characteristics ($\chi=2$).
Among all possible surfaces, we take the standard sphere as an example, see Figure 5. Therefore, direct integration of formula (4.14) and the usage of expression (5.1) imply the relation $\Lambda^{2}\mathcal{A}_{s}=4\pi\chi$, with $\mathcal{A}_{s}$ being the area of the sphere. For the sphere, the Euler characteristic is given by $\chi=2$. As a consequence, its area is fixed by $\mathcal{A}_{s}=8\pi/\Lambda^{2}$. And the radius $b_{s}$ of the sphere is thus determined by the simple formula
$b_{s}=\frac{\sqrt{2}}{\Lambda}\;.$ (5.8)
Substituting the numerical value of the cosmological constant, as computed in expression (3.5), we get
$b_{s}=15.7eV^{-1}\;.$ (5.9)
In SI units, the radius and the area of the sphere are $3.09\mu m$ and $119.98\mu m^{2}$, respectively.
Moreover, it is possible to relate the radii $b_{h}$ and $b_{s}$ by combining their area formulas. For instance, in the case of the open linearly connected $n$-holed torus, we have
$\mathcal{A}_{h}=(n-1)\mathcal{A}_{s}\;.$ (5.10)
Therefore, by considering as an example the system of $n$ identical torii of the previous Section, we get
$b_{h}=\sqrt{\frac{(n-1)}{2\pi n}}b_{s}\;.$ (5.11)
For the minimal ($n=2$) and extremal ($n\longrightarrow\infty$) cases we have a small range of values,
$\lim_{n\rightarrow 2}b_{h}$ $=$ $\frac{b_{s}}{2\sqrt{\pi}}\;,$
$\lim_{n\rightarrow\infty}b_{h}$ $=$ $\frac{b_{s}}{\sqrt{2\pi}}\;,$ (5.12)
in complete accordance with the first example in the previous Section.
6 Conclusions and perspectives
In this paper, we have shown that pure three-dimensional electrodynamics at the static limit corresponds to two-dimensional dilatonic gravity in Euclidean space. The Newtonian and cosmological constants are completely determined by the electric charge, as inferred in the Relations (3.3). Moreover, we were able to compute these constants, see Equation (3.5). Both of them are found to be several orders of magnitude larger than their four-dimensional experimental values, implying a kind of strong two-dimensional gravity.
We also demonstrated three possible solutions for such a gravity model in Sections 3 and 4, namely flat, hyperbolic, and spherical manifolds. These solutions essentially constitute the three possible situations of two-dimensional manifolds [63]. In Section 5, we employed the GB theorem to determine their sizes and some topological properties. Moreover, in all three solutions, the dilaton is a constant arbitrary field. A fact that is consistent with its electromagnetic origin as the scalar potential. The main properties of the solutions discussed in these Sections are described in what follows:
  • Flat manifolds: This solution is obtained from the action (3.4), originating directly from the mapping. In two dimensions, for several types of manifolds, the GB theorem, in formula (5.1), can be used to integrate the curvature scalar. Among these manifolds, since the curvature vanishes, we are restricted to manifolds with Euler characteristics zero. The possibilities we have inferred as examples are the cylinder (or, equivalently, a flat disk with a hole in it) and the flat torus (See Figure 1 for the examples). The flat torus is, of course, the only compact surface in these examples. Unfortunately, the size of these manifolds cannot be determined by the GB theorem.
  • Hyperbolic manifold: This solution is obtained by enforcing the shift constraint (4.3), related to the shift (4.2). The result is a JT-type of gravity with manifolds of constant negative curvature. Among hyperbolic surfaces, the tractroid, the catenoid, and the $n$-holed torus (See Figure 2 for the non-compact surfaces and Figures 3 and 4 for the compact examples) can be taken as examples. Since the catenoid and the tractroid are topologically equivalent to the cylinder, they have vanishing Euler characteristics, disabling the usage of the GB theorem (5.1) to determine the sizes of these surfaces.
    For the $n$-holed torii, on the other hand, we have applied the GB theorem in order to find their sizes in two examples: the open linearly connected $n$-holed torus (See Figure 3) and the closed linearly connected $n$-holed torus (Figure 4). The area of these manifolds depend on $n$ and $\Lambda$, as demonstrated in Expressions (5.3) and (5.7). Explicit computations were performed by considering a specific size of each torus, taken to be identical and characterized only by one parameter, the radius of the inner circle of the transversal cross-section of one torus. It was found an incredibly small radius for these manifolds. For the open $n$-holed torus with $n=2$, for example, the radius is just $0.87\mu m$. As $n$ increases, the radius asymptotically stabilizes at the value of $1.23\mu m$. For the closed $n$-holed torus, the radius does not depend on $n$ and is fixed to the value of $1.23\mu m$. Obviously, these values are computed from the value of the cosmological constant, calculated in (3.5), originally determined by the electric charge.
  • Spherical manifolds: This solution is obtained via the shift (4.11) and the (same as the hyperbolic solution) shift constraint (4.3). Constant positive curvature for JT gravity is obtained, namely (4.14). Therefore, the solution describes spherical manifolds. Among spherical surfaces, we pick as example the usual sphere (Figure 5), which is compact. Thence, GB theorem can easily be applied in order to determine the sphere radius in terms of the cosmological constant. Following the same steps developed in the $n$-holed torii examples, we were able to find that the radius of the sphere is astonishingly small, valuing $3.09\mu m$.
A more technical comment is that the constraint (4.3) (or the corresponding shifts of the curvature and spin-connection) is, in practice, equivalent to a choice of a background configuration. Particularly, hyperbolical or spherical ones. This background can only be chosen due to the presence of mass parameters in the action. Otherwise, a mass parameter would be required by hand in order to achieve JT gravity. Fortunately, we just do not have any mass parameter, but the specific parameter interpreted as a cosmological constant. Even if we did not have it, we recall that $G$ also carries mass dimension, allowing a different choice of curvature values.
Ultimately, the background can also be defined before the mapping. As discussed in Section 4, one could define a background field strength by the shift $F\longmapsto F\pm\Theta/2$ and the corresponding shift of the gauge potential $a$. The analogue of the shift constraint (4.3) would appear in terms of the fields of the MCS theory.
Another technical point is that the mapping (3.1) is fundamentally based on the isomorphism $U(1)\longmapsto SO(2)$. Therefore, we end up in a Euclidean gravity model. Minkowskian gravity could be obtained, in principle, by performing a Wick rotation at the gravity isometry group, $SO(2)\longmapsto SO(1,1)$, inducing a Minkowskian signature in the manifold metric. Thus, instead of hyperbolic and spherical manifolds, the surfaces would be Anti-de Sitter and de Sitter spacetimes, respectively. Nevertheless, the detailed study of the Wick rotation in the mapping we defined and the theories we studied is beyond the scope of the present work.
A remarkable feature of the model is that the relation between the gravitational constants and the electric charge, given in (3.3), indicates a possible duality because $G\sim 1/e^{2}$. Thence, if on one hand we have electrodynamics as a perturbative system, on the other hand we have a JT gravity model with a strong coupling parameter. Alternatively, there is no obstruction in starting from gravity with any value of $G$ and map it in static MCS theory while maintaining the relation between $G$ and $e$. In that case, one could describe standard JT gravity in terms of MCS static electromagnetism with some exotic value of electric charge. Thus, for example, a black hole configuration in JT gravity could be described by an electromagnetic system with a specific kind of charged fields. Such ideas also suggest that our model also has the potential to be employed in condensed matter systems, as discussed in the Introduction.
A few extra comments are in order. First, it is important to emphasize that we have used a map between pure static three-dimensional MCS electromagnetism and two-dimensional Euclidean gravity in the FOF. The electromagnetism side is a generic model with no characterized length scale (or energy scale). The only number we have considered is the fundamental charge of the electron $e$. On the gravity side, we found three possible solutions. Two of them have specific length scales of the order of $\mu m$ (micrometers). This feature originates from the fact that the curvatures of the exemplified surfaces are determined from the cosmological constant, while the latter is fixed by the electric charge. In essence, these two solutions are solutions of JT gravity with negative and positive curvatures.
In the same spirit, we can start in the gravity model we found and go back to electrodynamics at the static limit. Therefore, one could think of associating systems with characteristic lengths of the order of $\mu m$ with the properties of static MSC electrodynamics. Examples in Nature of objects and living beings of such small size are: suspended fine dust; the mitochondria; large viruses such as the Megavirus chilensis [75] and the Acanthamoeba polyphaga Mimivirus [76, 77]; some bacteria (e.g., the Escherichia coli, also known as the E. coli, [78]). Thus, perhaps, the surface of these systems could be mimicked by MCS static theory in flat space.
In truth, there are many perspectives for future investigations. We mention here some immediate possibilities. For instance, we can look at the four-dimensional case and its corresponding dimensional reduction (time and one space coordinate, for example). Naturally, one extra scalar field will emerge, and the dynamics can change. Another point is the inclusion of matter. Fermions could be considered in the same way we have considered them in [56], or one could simply include external source interactions. Of course, extra care will be demanded in the dimensional reduction of spinors [79, 80, 81]. It could also be interesting to study matter with different electric charges. In this case, the violation of the equivalence principle is expected since only one type of charge can be absorbed in the gauge field. Moreover, inclusion of matter may be relevant for black hole analysis [24, 26, 5]. In terms of time evolution, one can simply decompose space and time of three-dimensional electrodynamics and map to gravity only the space coordinates. The expected result would be that of surfaces and the dilaton evolving in time as a separate coordinate of the manifold. The same idea can be generalized to four dimensions with only one space coordinate dimensional reduction - in that case, two scalar fields will appear, possibly evolving in time together with the manifold.
Finally, it is worth mentioning that, in two dimensions, the map (3.1) is also consistent at the quantum level. Therefore, one can expect that the same feature is valid in the present case. Nevertheless, this study is beyond the scope of the present work.
Acknowledgements
The authors are in debt to J. Zanelli, M. S. Guimarães, and G. Sadovski for valuable discussions and ideas. This study was financed in part by The Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Brasil (CAPES) - Finance Code 001.